### v1

# Accelerating Diffusion Models in Offline RL via Reward-Aware Consistency Trajectory Distillation

Abstract
Although diffusion models have achieved strong results in decision-making tasks, their slow inference speed remains a key limitation. While the consistency model offers a potential solution, its applications to decision-making often struggle with suboptimal demonstrations or rely on complex concurrent training of multiple networks. In this work, we propose a novel approach to consistency distillation for offline reinforcement learning that directly incorporates reward optimization into the distillation process. Our method enables single-step generation while maintaining higher performance and simpler training. Empirical evaluations on the Gym MuJoCo benchmarks and long horizon planning demonstrate that our approach can achieve an $8.7\%$ improvement over previous state-of-the-art while offering up to 142×142$\times$142 × speedup over diffusion counterparts in inference time.
1 Introduction
Recent advances in diffusion models have demonstrated their remarkable capabilities across various domains (Song et al., 2020a; Karras et al., 2022; Liu et al., 2023; Chi et al., 2023; Janner et al., 2022), including decision-making tasks in reinforcement learning (RL). These models excel particularly in capturing multi-modal behavior patterns (Janner et al., 2022; Chi et al., 2023; Ajay et al., 2022) and achieving strong out-of-distribution generalization (Duan et al., 2025; Block et al., 2023), making them powerful tools for complex decision-making scenarios. However, their practical deployment faces a significant challenge: the computational overhead of the iterative sampling procedures, which requires numerous denoising steps to generate high-quality outputs.
To address this limitation, various diffusion acceleration techniques have been proposed, including ordinary or stochastic differential equations (ODE or SDE) solvers with flexible step sizes (Song et al., 2020a; Lu et al., 2022; Karras et al., 2022), sampling step distillation (Song et al., 2023; Kim et al., 2023) and improved noise schedules and parametrizations (Salimans and Ho, 2022; Song and Dhariwal, 2024). In particular, consistency distillation (Song et al., 2023) has emerged as one of the most promising solutions for image generation, in which a many-step diffusion model serves as a teacher to train a student consistency model that achieves comparable performance while enabling faster sampling through a single-step or few-step generation process.
This breakthrough has sparked considerable interest in applying consistency distillation to decision-making tasks. However, current applications either adopt a behavior cloning approach (Lu et al., 2024; Prasad et al., 2024; Wang et al., 2024) or integrate few-step diffusion based samplers in actor-critic frameworks (Chen et al., 2023; Ding and Jin, 2023; Li et al., 2024). While promising, these approaches face inherent challenges: behavior cloning performs well only with expert demonstrations but struggles with suboptimal data (e.g., median-quality replay buffers), while the actor-critic paradigm introduces concurrent training of multiple networks and sensitive hyperparameters, exacerbating training complexity, instability, and computational overhead.
This raises an important question: can we develop a more effective approach to consistency distillation specifically tailored for offline RL? We address this challenge by introducing a novel method that directly incorporates reward optimization into the consistency distillation process. Our approach begins with a pre-trained unconditional diffusion policy and augments the standard consistency trajectory distillation (Kim et al., 2023) with an explicit reward objective. The vanilla consistency trajectory distillation helps the student model cover the diverse behavior patterns learned by the teacher. The additional objective encourages the student consistency model to sample actions that yield higher rewards, effectively steering the model toward selecting optimal trajectories from the multi-modal distributions captured by the teacher diffusion model.
Another key advantage of our method lies in its training simplicity. The reward model can be trained independently from the teacher diffusion model and the distillation process, avoiding the complexity of concurrent multi-network training present in actor-critic methods. Our method also eliminates the need for training noise-aware reward models, unlike existing guided diffusion sampling approaches (Janner et al., 2022). By integrating reward optimization directly into the distillation process, our method achieves superior performance, enables efficient single-step generation, and maintains straightforward training procedures.
We demonstrate the performance and sampling efficiency of our RACTD on the suboptimal heterogenous D4RL Gym-MuJoCo benchmark and challenging long-horizon planning task Maze2d (Fu et al., 2020). Our method demonstrates both superior performance and substantial sampling efficiency compared to existing approaches, achieving an $8.7\%$ improvement compared to existing state-of-the-art (SOTA) and a $142$-fold reduction in sampling time.
Our contributions include: (1) We propose a novel reward-aware consistency distillation method for offline RL that enables single-step generation while achieving superior performance, (2) We demonstrate that our approach enables decoupled training without the complexity of concurrent multi-network optimization or noise aware reward model training, (3) Through comprehensive experiments on multi-modal suboptimal dataset and long-horizon planning tasks, we show that our method achieves $8.7\%$ improvement over prior SOTA while achieving up to 142×142$\times$142 × speedup. We will publicly release our code upon acceptance of this paper.
Figure 1: Overview of Reward Aware Consistency Trajectory Distillation (RACTD). We incorporate reward guidance with consistency trajectory distillation to train a student model that can generate actions with high rewards with only one denoising step.
2 Background
2.1 Problem Setting
In this paper we consider the classic setting of offline reinforcement learning, where the goal is to learn a policy $\pi$ to generate actions that maximize the expected cumulative discounted reward in a Markov decision process (MDP). A MDP is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},R,\gamma)$, where $\mathcal{S}$ is the set of possible states $\bm{s}\in\mathcal{S}$, $\mathcal{A}$ is the set of actions $\bm{a}\in\mathcal{A}$, $\mathcal{P}(\bm{s}^{\prime}\mid\bm{s},\bm{a})$ is the transition dynamics, $R(\bm{s},\bm{a})$ is a reward function, and $\gamma\in[0,1]$ is a discount factor. In offline RL, we further assume that the agent can no longer interact with the environment and is restricted to learning from a size $M$ static dataset $\mathcal{D}=\{\bm{\tau}_{i}\}_{i=1}^{M}$, where $\bm{\tau}=(\bm{s}_{0},\bm{a}_{0},r_{0},\bm{s}_{1},\bm{a}_{1},r_{1},\dots,\bm{s}_{H},\bm{a}_{H},r_{H})$ represents a rollout of episode horizon $H$ collected by following a behavior policy $\pi_{\beta}$.
Mathematically, we want to find a policy $\pi^{*}$ that
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\bm{\tau}\sim\pi}\Bigl{[}\sum_{n=0}^{H}\gamma^{n}\,R\bigl{(}\bm{s}_{n},\bm{a}_{n}\bigr{)}\Bigr{]}$ (1)
subject to the constraint that all policy evaluation and improvement must rely on $\mathcal{D}$ alone.
2.2 Diffusion Models
Diffusion models generate data by learning to reverse a gradual noise corruption process applied to training examples. Given a clean data sample $\bm{x}_{0}$, we define $\bm{x}_{t}$ for $t\in[0,T]$ as increasingly noisy versions of $\bm{x}_{0}$. The forward (or noising) process is commonly formulated as an Itô stochastic differential equation (SDE):
$\mathrm{d}\bm{x}=f(\bm{x},t)\mathrm{d}t+g(t)\mathrm{d}w$ (2)
where $w$ is a standard Wiener process. As $t$ approaches the final timestep $T$, the distribution of $\bm{x}_{T}$ converges to a known prior distribution, typically Gaussian. At inference time, the model reverses this corruption process by following the corresponding reverse-time SDE, which depends on the score function $\nabla_{\bm{x}}\log p_{t}(\bm{x})$. In practice, this score function is approximated by a denoiser network $D_{\phi}$, enabling iterative denoising from $\bm{x}_{T}$ back to $\bm{x}_{0}$. An alternative, deterministic interpretation of the reverse process is given by the probability flow ODE (PFODE):
$\mathrm{d}\bm{x}=\Bigl{[}f(\bm{x},t)-\tfrac{1}{2}g(t)^{2}\nabla_{\bm{x}}\log p_{t}(\bm{x})\Bigr{]}\mathrm{d}t$ (3)
which preserves the same marginal distribution $p_{t}(\bm{x})$ as the reverse SDE at each timestep $t$. This ODE formulation often enables more efficient sampling through larger or adaptive step sizes without significantly compromising the sample quality.
EDM (Karras et al., 2022) refine both the forward and reverse processes through improved noise parameterization and training objectives. Concretely, they reparametrize the denoising score matching (DSM) loss so that the denoiser network learns to predict a scaled version of the clean data:
$\mathcal{L}_{\mathrm{EDM}}={\mathbb{E}}_{t,\bm{x}_{0},\bm{x}_{t}|\bm{x}_{0}}\left[d(\bm{x}_{0},D_{\phi}(\bm{x}_{t},t))\right]$ (4)
where $d$ is a distance metric in the clean data space. In this paper, we train an EDM model as the teacher using the pseudo huber loss as $d$ following Prasad et al. (2024). At inference time, EDM solves the associated PFODE with a 2nd-order Heun solver.
2.3 Consistency Trajectory Distillation
The iterative nature of the diffusion sampling process introduces significant computational overhead. Among various acceleration techniques proposed, consistency distillation (Song et al., 2023) has emerged as a particularly effective approach. The core idea is to train a student model that can emulate the many-step denoising process of a teacher diffusion model in a single step.
Building upon this framework, Kim et al. (2023) introduced Consistency Trajectory Models (CTM). Instead of learning only the end-to-end mapping from noise to clean samples, CTM learns to predict across arbitrary time intervals in the diffusion process. Specifically, given three arbitrary timesteps $0\leq k<u<t\leq T$, CTM aims to align two different paths to predict $\bm{x}_{k}$: (1) direct prediction from time $t$ to $k$ using the student model, and (2) a two-stage prediction that first uses a numerical solver (e.g., Heun) with the teacher model to predict from time $t$ to $u$, and then uses the student model to predict from time $u$ to $k$.
Since the distance metric $d$ is defined on the clean data space and may not be well-defined in the noisy data space, in practice we further map all predictions to time 00 using the student model. Formally, denote the student model as $G_{\theta}$, the CTM loss is defined as:
$\small\mathcal{L}_{\mathrm{CTM}}=\mathbb{E}\left[d\left(G_{sg(\theta)}(\hat{\bm{x}}_{k}^{(t)},k,0),G_{sg(\theta)}(\bm{x}_{k}^{(t,u)},k,0)\right)\right]$ (5)
where $G_{\theta}(\bm{x}_{t},t,u)$ represents the student prediction from time $t$ to $u$ given noisy sample $\bm{x}_{t}$ at time $t$, $\mathrm{sg}(\theta)$ represents stop-gradient student parameters and
$\hat{\bm{x}}_{k}^{(t)}=G_{\theta}(\bm{x}_{t},t,k),$ $\bm{x}_{k}^{(t,u)}=G_{\mathrm{sg}(\theta)}(\mathrm{Solver}(\bm{x}_{t},t,u;\phi),u,k)$ (6)
Here $\mathrm{Solver}(\bm{x}_{t},t,u;\phi)$ is the numerical solver result from time $t$ to $u$ using the teacher model $D_{\phi}$ given noisy sample $\bm{x}_{t}$ at time $t$.
In addition to the CTM loss, consistency trajectory distillation also incorporates the DSM loss to enforce its samples to be close to the training data. The DSM loss for the student model is the same as the one for EDM in Equation 4:
$\mathcal{L}_{\mathrm{DSM}}={\mathbb{E}}_{t,\bm{x}_{0},\bm{x}_{t}|\bm{x}_{0}}\left[d(\bm{x}_{0},G_{\theta}(\bm{x}_{t},t,0))\right]$ (7)
In Figure 2 we provide a visualization of these objectives on a PFODE trajectory. After the distillation, the student model can perform “anytime-to-anytime” jumps along the PFODE trajectory. One-step sampling can then be achieved by calculating $\hat{\bm{x}}^{(T)}_{0}=G_{\theta}(\bm{x}_{T},T,0)$.
Figure 2: Visualization of CTM loss, DSM loss and reward loss.
3 Method
3.1 Motivation and Intuition
Diffusion models and their consistency-based counterparts have demonstrated promising results in decision-making tasks, particularly in capturing multimodal behavior patterns (Janner et al., 2022; Chi et al., 2023; Ajay et al., 2022). Common recipes for using these models in decision-making generally fall into one of the three paradigms: (1) training a diffusion or consistency model via behavior cloning and deploying it directly as a policy (Chi et al., 2023; Ajay et al., 2022); (2) integrating the model into actor-critic frameworks (Wang et al., 2022; Hansen-Estruch et al., 2023; Ding and Jin, 2023); or (3) using the model as a planner through guided diffusion sampling (Janner et al., 2022).
While behavior cloning pipelines perform well when trained on expert demonstrations, they often struggles with suboptimal datasets (e.g., medium-replay buffers) collected from diverse behavior policies. Such datasets typically exhibit complex multimodal behavior patterns, where only some modes lead to high rewards. Although one could potentially use rejection sampling to filter out low-reward training data, this approach becomes prohibitively sample inefficient, particularly as the quality of the training data deteriorates. On the other hand, to generate high reward actions, actor-critic approaches require concurrent training of multiple neural networks with sensitive hyperparameters. Finally, guided diffusion sampling necessitates training noise-aware reward models and multi-step sampling, which could be detrimental for time-sensitive decision-making tasks like self-driving.
So how can we better leverage potentially suboptimal datasets to design a diffusion-based single-step sampling model with simple training procedure? Our key idea is to utilize the multimodal information captured by the teacher model and encourage the student model to favor the high reward modes. We achieve this by incorporating a reward objective directly in the consistency distillation process. Since our student model can achieve single-step denoising, we can incorporate a reward model trained in the clean sample space and avoid the multi-step reward optimization for diffusion models.
3.2 Modeling Action Sequences
Before introducing our reward-aware consistency trajectory distillation, we would like to first clearly establish the modeling formulation in our method. When applying diffusion models to RL, several modeling choices are available: modeling actions (as a policy), modeling rollouts (as a planner), or modeling state transitions (as a world model). While more comprehensive modeling approaches can offer advantages, particularly in long-horizon tasks, they also introduce additional complexity and computational challenges.
Following Chi et al. (2023), we adopt a balanced approach: modeling a fixed-length sequence of future actions conditioned on a fixed-length sequence of observed states. This formulation ensures consecutive actions form coherent sequences, and reduces the chances of generating idle actions.
[…]
During execution, we can either execute only the first predicted action $\bm{a}_{n}$ in the environment before replanning at the next step, or follow the entire predicted sequence of actions at once.
3.3 Reward-Aware Consistency Trajectory Distillation
Having established the formulation, we now present our approach to integrating reward optimization into the consistency trajectory distillation process.
Let $R_{\psi}$ be a pre-trained differentiable return-to-go network (i.e. reward model) that takes the state $\bm{s}_{n}$ and action $\bm{a}_{n}$ at rollout time $n$ as inputs and predicts the future discounted cumulative reward $\hat{r}_{n}=\sum_{j=0}^{H-n}\gamma^{j}r_{n+j}$. When the student model generates a prediction $\vec{\bm{a}}_{n}\mid\vec{\bm{s}}_{n}=\hat{\bm{x}}^{(T)}_{0}=G_{\theta}(\bm{x}_{T},T,0)$, we extract the action at time $n$, denoted as $\hat{\bm{a}}_{n}$, from the predicted sequence and pass it along with $\bm{s}_{n}$ to the frozen reward model $R_{\psi}$ to estimate $\hat{r}_{n}$. The goal of our reward-aware training is to maximize the estimated discounted cumulative reward. Mathematically, the reward objective is defined as
$\mathcal{L}_{\mathrm{Reward}}=-R_{\psi}(\vec{\bm{s}}_{n},\hat{\bm{a}}_{n})$ (8)
The final loss for reward-aware consistency trajectory distillation (RACTD) combines all three terms:
$\mathcal{L}=\alpha\mathcal{L}_{\mathrm{CTM}}+\beta\mathcal{L}_{\mathrm{DSM}}+\sigma\mathcal{L}_{\mathrm{Reward}}$ (9)
where $\alpha$, $\beta$, and $\sigma$ are hyperparameters to balance different loss terms.
3.4 Decoupled Training
A key advantage of our method of combining reinforcement learning, diffusion models, and consistency distillation is the ability to support fully decoupled training of all components. Traditional actor-critic frameworks, which are commonly used to incorporate diffusion models into reinforcement learning, require simultaneous training of multiple neural networks. This concurrent optimization presents considerable challenges, often demanding extensive hyperparameter tuning and careful balancing of different learning objectives.
Guided diffusion sampling, as proposed in Janner et al. (2022), offers an alternative approach by taking inspiration from classifier guided diffusion (Dhariwal and Nichol, 2021; Song et al., 2020b). However, these classifiers (i.e. reward models) require noise-aware training that cannot be separated from the diffusion model. Also, predicting the correct reward from highly corrupted input could be very challenging, which can lead to inaccurate guidance that accumulates during its multi-step sampling process.
Our method, in contrast, fully leverages the advantages of single-step denoising models by operating entirely in the noise-free state-action space. This design choice enables the reward model to provide stable and effective signals without requiring noise-aware training. Importantly, the reward model can be pre-trained completely decoupled from the teacher model and distillation process. This separation not only simplifies the training process but also allows for flexible integration of different reward models using the same teacher model.
3.5 Reward Objective as Mode Selection
In offline RL, models often have to learn from datasets containing behaviors of varying quality. While diffusion models excel at capturing these diverse behavioral modes, they inherently lack the ability to differentiate between actions that lead to high versus low rewards. Our RACTD addresses this limitation by transforming the reward-agnostic teacher diffusion sampling distribution into one that preferentially samples from high-reward modes, We empirically verify this through a comparative analysis using the D4RL hopper-medium-expert dataset Fu et al. (2020), which contains an equal mixture of expert demonstrations and suboptimal rollouts from a partially trained policy.
Figure 3: The reward distribution of the D4RL hopper-medium-expert dataset and 100 rollouts from an unconditioned teacher, an unconditioned student, and RACTD.
Figure 3 illustrates the reward distributions of rollouts sampled from three models: the unconditioned teacher, unconditioned student, and RACTD. The dataset (grey) exhibits two distinct modes corresponding to medium-quality and expert rollouts. The unconditioned teacher model (blue) accurately captures this bimodal distribution, and the unconditioned student model (orange) faithfully replicates it. In contrast, our RACTD (green) concentrates its samples on the higher-reward mode, demonstrating that our reward guidance effectively identifies and selects optimal behaviors from the teacher’s multi-modal distribution. We also include the discussion between sample diversity and mode selection in Appendix D.
4 Experiment
In this section, we conduct experiments to demonstrate: (1) the effectness of our RACTD, especially when the teacher model is trained with suboptimal data, (2) the capability of accurately capturing the complex and diverse high dimensional behavior patterns with only one denoising step, and (3) the speed-up achieved over the teacher model and existing policy-based diffusion models.
4.1 Offline RL
Baselines
We compare our approach against a comprehensive set of baselines, including vanilla behavior cloning (BC) and Consistency Policy (Consistency BC) (Ding and Jin, 2023); model-free RL algorithms CQL (Kumar et al., 2020) and IQL (Kostrikov et al., 2021); model-based algorithms Trajectory Transformer (TT) (Janner et al., 2021), MOPO (Yu et al., 2020), MOReL (Kidambi et al., 2020), MBOP (Argenson and Dulac-Arnold, 2020); autoregressive model Decision Transformer (DT) Chen et al. (2021); diffusion-based planner Diffuser Janner et al. (2022); and diffusion-based actor-critic methods Diffusion QL (Wang et al., 2022) and Consistency AC (Ding and Jin, 2023).
Setup
We first evaluate our method and the baselines on D4RL Gym-MuJoCo (Fu et al., 2020), which is a popular offline RL benchmark that contains three environments (hopper, walker2d, halfcheetah) of mixtures of varying quality data (medium-replay, medium, medium-expert).
Following the setups in the prior works, we employ online model selection for evaluation by using the best-performing checkpoint observed during training. Results for non-diffusion-based models and Diffuser are taken from Janner et al. (2022), and results for Diffusion QL and Consistency AC/BC are sourced from their respective papers. The results for RACTD are reported as the mean and standard error over 100 planning seeds. We use $h=1,c=16$ and closed-loop plannign in all experiments.
Table 1: (Offline RL) Performance of RACTD and a variety of baselines on the D4RL Gym-MujoCo benchmark. The best score is emphasized in bold and the second-best is underlined.
Method Medium Expert Medium Medium Replay   Avg   NFE
HalfCheetah Hopper Walker2d HalfCheetah Hopper Walker2d HalfCheetah Hopper Walker2d
BC $55.2$ $52.5$ $107.5$ $42.6$ $52.9$ $75.3$ $36.6$ $18.1$ $26.0$ $\ 51.9$
CQL $91.6$ $105.4$ $108.8$ $44.0$ $58.5$ $72.5$ $45.5$ $95.0$ $77.2$ $\ 77.6$
IQL $86.7$ $91.5$ $109.6$ $47.4$ $66.3$ $78.3$ $44.2$ $94.7$ $73.9$ $\ 77.0$
DT $86.8$ $107.6$ $108.1$ $42.6$ $67.6$ $74.0$ $36.6$ $82.7$ $66.6$ $\ 74.7$
TT $95.0$ $110.0$ $101.9$ $46.9$ $61.1$ $79.0$ $41.9$ $91.5$ $82.6$ $\ 78.9$
MOPO $63.3$ $23.7$ $44.6$ $42.3$ $28.0$ $17.8$ $53.1$ $67.5$ $39.0$ $\ 42.1$
MOReL $53.3$ $108.7$ $95.6$ $42.1$ $95.4$ $77.8$ $40.2$ $93.6$ $49.8$ $\ 72.9$
MBOP 105.9 $55.1$ $70.2$ $44.6$ $48.8$ $41.0$ $42.3$ $12.4$ $9.7$ $\ 47.8$
Diffusion QL $\underline{97.2}$$\pm 0.4$ $\underline{112.3}$$\pm 0.8$ $111.2$$\pm 0.9$ $51.5$$\pm 0.3$ $96.6$$\pm 3.4$ $\underline{87.3}$$\pm 0.5$ $48.3$$\pm 0.2$ $\underline{102.0}$$\pm 0.4$ 98.0 ±$\pm 0.5$ 89.389.3\ 89.389.3 $5$
Consistency AC $89.2$$\pm 3.3$ $106.0$$\pm 1.3$ $\underline{111.6}$$\pm 0.7$ 71.9 ±$\pm 0.8$ $\underline{99.7}$$\pm 2.3$ $84.1$$\pm 0.3$ 62.7 ±$\pm 0.6$ $100.4$±$\pm 0.6$ $83.0$$\pm 1.5$ $\underline{89.8}$ $2$
Consistency BC $39.6$$\pm 3.4$ $96.8$$\pm 4.6$ $\underline{111.6}$$\pm 0.7$ $46.2$$\pm 0.4$ $78.3$$\pm 2.6$ $84.1$$\pm 0.3$ $45.4$$\pm 0.7$ $100.4$$\pm 0.6$ $80.8$$\pm 3.4$ 75.975.9\ 75.975.9 $2$
Diffuser $88.9$$\pm 0.3$ $103.3$$\pm 1.3$ $106.9$$\pm 0.2$ $42.8$$\pm 0.3$ $74.3$$\pm 1.4$ $79.6$$\pm 0.55$ $37.7$$\pm 0.5$ $93.6$$\pm 0.4$ $70.6$$\pm 1.6$ 77.577.5\ 77.577.5 $20$
RACTD(Ours) $95.9$$\pm 1.5$ 129.0 ±$\pm 1.3$ 119.6 ±$\pm 1.9$ $\underline{59.3}$$\pm 0.2$ 121.0 ±$\pm 0.5$ 98.9 ±$\pm 1.6$ $\underline{57.9}$$\pm 1.0$ 104.9 ±$\pm 2.1$ $\underline{91.5}$$\pm 2.6$ 97.6 1
Results
As shown in Table 1, RACTD achieves the best or second best performance in almost all task, and outperforms the best baseline in overall average score by a substantial margin. It consistently outperforms the diffusion-based planner Diffuser in all environments, and exceeds the performance of consistency-based actor-critic baseline Consistency AC in all hopper and walker2d tasks.
4.2 Long Horizon Planning
Next, we showcase the effectiveness of RACTD in complex high-dimensional long-horizon planning and its flexibility to adapt to goal-conditioned tasks. Previously, Diffuser has shown great potential in open-loop long-horizon planning, but requires a significantly larger number of denoising steps compared to closed-loop planning like MuJoCo. We demonstrate that our model can achieve superior performance with a single-step denoising process under the same problem formulation.
Setup
We test this ability on D4RL Maze2d (Fu et al., 2020), which is a sparse reward long-horizon planning task where an agent may take hundreds of steps to reach the goal in static environments. Following the setup in Janner et al. (2022), we use a planning horizon $128,256,384$ for U-Maze, Medium and Large respectively. We perform open-loop planning by generating the entire state sequence followed by a reverse dynamics model to infer all the actions from the predicted state sequence. The reward model returns $1$ if the current state reaches the goal and 00 otherwise. The baseline results are reported from Janner et al. (2022) and RACTD results are reported as the mean and standard error of 100 planning seeds.
Results
As shown in Table LABEL:tab:long, both Diffuser and RACTD outperform previous model-free RL algorithms CQL and IQL, and MPPI which usese ground-truth dynamics. Our approach surpasses the diffusion baseline in almost all settings, highlighting its ability to effectively capture complex behavioral patterns and high-dimensional information in the training dataset. Notably, the planning dimension for this task (384 for the Large Maze) is substantially higher than that of MuJoCo tasks (16). As a result, Diffuser requires significantly more denoising steps (256 for the Large Maze) compared to MuJoCo (20 steps). On the contrary, despite the increased task complexity, RACTD still only requires a single denoising step to achieve 11.6×11.6$\times$11.6 × performance boost.
Table 2: (Long-horizon planning) The performance of RACTD, Diffuser, and prior model-free algorithms in the Maze2D environment. The best score is in bold and the second-best is underlined.
Method U-Maze Medium Large Avg Score
Score NFE Time (s) Score NFE Time (s) Score NFE Time (s)
MPPI $33.2$ $10.2$ $5.1$ $16.2$
CQL $5.7$ $5$ $12.5$ $7.7$
IQL $47.4$ $34.9$ $58.6$ $47.0$
Diffuser $113.9$$\pm 3.1$ $64$ $1.664$ 121.5 ±$\pm 2.7$ $256$ $4.312$ $123.0$$\pm 6.4$ $256$ $5.568$ $119.5$
CTD $123.4$$\pm 1.0$ 1 0.029 $119.8$$\pm 4.1$ 1 0.047 $127.1$$\pm 6.4$ 1 0.049 $123.4$
RACTD (Ours) 125.7 ±$\pm 0.6$ 1 0.029 130.8 ±$\pm 1.8$ 1 0.047 143.8 ±$\pm 0.0$ 1 0.049 133.4
4.3 Inference Time Comparison
Beyond performance improvements, a key contribution of our work is significantly accelerating diffusion-based models for decision-making tasks. The primary computational bottleneck in diffusion models arises from the multiple function evaluations (NFEs) required by the iterative denoising process. By reducing the number of denoising steps to a single NFE, our approach achieves a speed-up roughly proportional to the number of denoising originally required.
Setup
To evaluate sampling efficiency, we compare RACTD with different samplers, including DDPM (Ho et al., 2020), DDIM (Song et al., 2020a), and EDM (Karras et al., 2022) using the same network architecture as our teacher model. Additionally, we report the efficiency of Diffuser. Note that since Diffuser employs a different model architecture and generates future state-action pairs, its sampling time may also be influenced by these factors. Table 4 and Table LABEL:tab:long present the wall clock sampling time and NFE for MuJoCo (hopper-medium-replay) and Maze2d. All experiments are conducted on one NVIDIA Tesla V100-SXM2-32GB.
Results
In hopper-medium-replay, RACTD achieved 20×20$\times$20 × reduction in NFEs and a 43×43$\times$43 × speed-up compared to Diffuser. Additionally, our student model requires 80×80$\times$80 × fewer NFEs and samples 142×142$\times$142 × faster than the teacher model. In Maze2d, RACTD significantly accelerates computation compared to Diffuser, achieving approximately 57×57$\times$57 ×, 92×92$\times$92 ×, and 114×114$\times$114 × speed-ups on Umaze, Medium and Large mazes, by reducing NFEs by a factor of $256$ for the Medium and Large mazes. Furthermore, we observe that more denoising steps lead to better model performance. This highlights the advantage of distilling from a slow but high-performing teacher model, which enables better performance compared to training a diffusion model directly with fewer denoising steps.
5 Ablation Study
We ablate over (1) the impact of the reward objective in both student and teacher model training, (2) the effect of different reward loss weights on model performance and training stability, and (3) the impact of increasing denoising steps with our student model.
Table 3: Wall clock time and NFEs per action for different samplers and Diffuser on MuJoCo hopper-medium-replay.
Method Time (s) NFE Score
Diffuser $0.644$ $20$ $93.6$
DDPM $0.236$ $15$ $24.2$
DDIM $0.208$ $15$ $60.6$
EDM (Teacher) $2.134$ $80$ 114.2
RACTD (Ours, Student) 0.015 1 $104.9$
Figure 4: Ablation on reward objective weight.
5.1 Impact of Reward Objective
To understand the unique advantages of incorporating reward objectives during distillation, we conduct a systematic comparison across four model configurations: the baseline combination of an unconditioned teacher and student, a reward-aware teacher paired with an unconditioned student, a fully reward-aware teacher-student pair, and our proposed RACTD, which combines an unconditioned teacher with a reward-aware student. The results for hopper-medium-replay and walker-medium is shown in Table 5 and Table 10.
Our analysis reveals that while incorporating reward objectives at any stage yields substantial improvements, optimal performance is achieved through our RACTD framework, which combines an unconditioned teacher with reward-aware student distillation. Our method allows the teacher to capture a comprehensive range of behavioral patterns, while enabling the student to selectively distill the most effective strategies. Although incorporating reward objectives in the teacher model also enhances performance, this approach risks discarding suboptimal behaviors that may be valuable in novel testing scenarios, potentially limiting the model’s generalization capabilities.
Table 4: We compare incorporating the reward model in different stages of training on MuJoCo hopper-medium-replay. Results are presented as the mean and standard error across 100 seeds.
Hopper Medium-Replay Unconditioned teacher Reward-Aware teacher
Unconditioned student $50.8$$\pm 0.3$ $96.2$$\pm 0.2$
Reward-Aware student 104.9 ±$\pm 2.1$ $96.0$$\pm 0.3$
Table 5: Inference time, NFE, and score comparison for student model multi-step sampling on MuJoCo hopper-medium-replay.
NFE Time(s) Score
1 0.0147 $104.9$$\pm 2.1$
2 $0.0241$ 109.8 ±$\pm 0.9$
3 $0.0377$ $108.7$$\pm 0.1$
4 $0.0517$ $107.9$$\pm 1.6$
5.2 Effect of Reward Objective Weights
The reward loss weight is a crucial hyperparameter in our pipeline that impacts both training stability and performance. We plot the reward curve achieved over 200 epochs of student training with reward loss weights ranging from $[0.3,0.5,0.7,1.5]$ on MuJoCo hopper-medium-replay in Figure 4. The mean and standard error are reported across 20 rollouts from intermediate checkpoints.
With lower weights, increasing the weight leads to higher performance and relatively stable training. However, when the weight is too high (e.g. $1.5$ in this plot), evaluation initially increases but fluctuates as training progresses. This occurs when the reward loss dominates DSM and CTM losses, resulting in unstable training.
5.3 Number of Sampling Steps
Since our student model is trained for anytime-to-anytime jumps, it naturally extends to multi-step denoising without additional training. Following the approach in Song et al. (2023), given intermediate denoising timesteps $0<t_{1}<t_{2}<T$, we first denoise from $T$ to 00 as usual. We then add noise again to $t_{1}$ and denoise it back to 00, and repeat this process for $t_{2}$. This iterative refinement can enhance generation quality. We evaluate the student using 2, 3, and 4 denoising steps as reported in Table 5. While multi-step sampling improves performance, we observe that gains do not scale linearly with the number of denoising steps.
6 Related Work
Diffusion Models in Reinforcement Learning
Diffusion models have emerged as a powerful approach for decision-making tasks in reinforcement learning Janner et al. (2022); Ajay et al. (2022); Wang et al. (2022); Hansen-Estruch et al. (2023); Chi et al. (2023). The integration of diffusion models into RL frameworks typically follows two main paradigms: actor-critic approaches, where diffusion models serve as policy or value networks (Wang et al., 2022; Hansen-Estruch et al., 2023), and policy-based approaches, where diffusion models directly generate action trajectories Janner et al. (2022); Ajay et al. (2022); Chi et al. (2023). While these methods have demonstrated impressive performance on standard RL benchmarks, their practical deployment is hindered by the slow sampling time inherent to vanilla diffusion policies based on DDPM Ho et al. (2020). This limitation poses particular challenges for real-world applications such as robotics, where speed is crucial.
Accelerating Diffusion Model Sampling
Various approaches have been proposed to accelerate the sampling process in diffusion models. One prominent direction leverages advanced ODE solvers to reduce the number of required denoising steps (Song et al., 2020a; Karras et al., 2022; Lu et al., 2022). Another line of work explores knowledge distillation techniques Luhman and Luhman (2021); Salimans and Ho (2022); Berthelot et al. (2023); Kim et al. (2023); Song et al. (2023), where student models learn to take larger steps along the ODE trajectory. Particularly, consistency models Song et al. (2023) leverage the PFODE to learn direct mappings from noise to sample space. Consistency trajectory models (Kim et al., 2023) enable one-step sampling by learning anytime-to-anytime jumps along the PFODE trajectory, achieving performance that surpasses teacher models.
Consistency Models in Decision Making
Consistency models have emerged as a promising policy class for behavior cloning from expert demonstrations in robotics (Lu et al., 2024; Prasad et al., 2024; Wang et al., 2024). In RL, several works have enhanced actor-critic methods by replacing traditional diffusion-based value/policy networks with consistency models, showing faster inference and training speed (Chen et al., 2023; Ding and Jin, 2023; Li et al., 2024). These approaches directly incorporate consistency loss (Song et al., 2023) into the value/policy network training, rather than distilling a separate student model. While Wang et al. made some initial attempts to apply consistency distillation in policy-based RL through classifier-free guidance and reverse dynamics, their approach requires two NFEs and under-performs the state-of-the-art. In contrast, RACTD is a straightforward approach of using a separate reward model and incorporating reward objective during student distillation, achieving superior performance with only one NFE.
7 Conclusion
In this work, we address the challenge of diffusion policy acceleration by introducing reward-aware consistency trajectory distillation (RACTD), which predicts high-reward actions in a single denoising step. RACTD uses a pre-trained diffusion teacher model and a separately trained reward model, leveraging the teacher’s ability to capture multi-modal distributions while prioritizing higher-reward modes to generate high-quality samples from suboptimal training data. Its decoupled training approach avoids the complex concurrent optimization of multiple networks and enables the use of a standalone, noise-free reward model. RACTD outperforms previous state-of-the-art by $8.7\%$ while accelerating its diffusion counterparts up to a factor of $142$.
Acknowledgment
This work was partly supported by DSTA, ONR N000142312368 and ONR MURI N00014-25-1-2116. FT was partially supported by the U.S. Army Futures Command under Contract No. W519TC-23-C-0030 during the project.
Appendix A Model Architecture
We follow the model architecture used in Chi et al. [2023] and Prasad et al. [2024] and continue to use the 1D temporal CNN layer for our Unet and FiLM layers to process the conditioning information.
A.1 Model Sizes for Maze2d
Model parameters for teacher models of Umaze, Medium, and Large Maze are shown below in Table 6. The student model has the same architecture as the teacher model except it also takes one extra variable of denoising timestep as conditioning.
Parameter Umaze Medium Large
diffusion_step_embed_dim 256 256 256
down dims [256, 512, 1024] [512, 1024, 2048] [256, 512, 1024, 2048]
horizon 128 256 384
kernel size 5 5 5
Table 6: Model parameters for Unet in Maze2d.
A.2 Model Sizes for Gym-MuJoCo
Unet parameters for teacher and student models in the MuJoCo task are shown below in Table 7. Model sizes are fixed through 9 different environments.
Parameter MuJoCo
diffusion_step_embed_dim 128
down dims [512, 1024, 2048]
horizon 16
kernel size 5
Table 7: Model parameters for Unet in MuJoCo.
We follow the setup in Janner et al. [2022], where we use Linear layers and Mish layers Misra [2019] for the reward model. Reward model architecture remains the same across all MuJoCo benchmarks. Model parameters are shown below in Table 8.
Parameter MuJoCo
layer dimensions [32, 64, 128, 256]
Table 8: Reward model parameters in MuJoCo.
Appendix B Training Details
Our models are trained on D4RL dataset [Fu et al., 2020], which was released under Apache-2.0 license.
B.1 Noise Scheduler
We follow the setup in Prasad et al. [2024] and use EDM noise scheduler [Karras et al., 2022] for the teacher model. Particularly, discretization bins are chosen to be 80.
Student model used CTM scheduler [Kim et al., 2023] also with discretization bins of 80.
B.2 Reward Discount Factor
We use $\gamma=0.997$ for hopper, $\gamma=0.99$ for walker2d, and $\gamma=0.9$ for halfcheetah, along with an early termination penalty of -100.
B.3 Weight of Different Losses
The weights for CTM, DSM, and Reward loss we used in the experiment are shown below in Table 9. Generally, if the training dataset includes more expert samples, the weight for reward guidance is smaller. A reward weight of $0.0$ resembles behavior cloning with consistency trajectory distillation. We found that as long as the loss weights are chosen to keep the individual loss terms within the same order of magnitude, the model will achieve reasonable performance.
Parameter CTM DSM Reward
hopper-medium-replay 1.0 1.0 1.0
hopper-medium 1.0 1.0 3.0
hopper-medium-expert 1.0 1.0 0.0
walker2d-medium-replay 1.0 1.0 1.0
walker2d-medium 1.0 1.0 0.4
walker2d-medium-expert 1.0 1.0 0.2
halfcheetah-medium-replay 1.0 1.0 1.0
halfcheetah-medium 1.0 1.0 0.5
halfcheetah-medium-expert 1.0 1.0 0.0
Table 9: Weights for CTM, DSM, and Reward loss used in MuJoCo benchmark.
Appendix C More Ablations
C.1 walker-medium
Table 10: We compare incorporating the reward model in different stages of training on MuJoCo walker-medium. Results are presented as the mean and standard error across 100 seeds.
Walker Medium Unconditioned teacher Reward-Aware teacher
Unconditioned student $93.3$$\pm 1.8$ $97.0$$\pm 1.0$
Reward-Aware student 98.9 ±$\pm 1.6$ $94.5$$\pm 2.6$
C.2 Comparing with fast sampling algorithms for Maze2d
Table 11: We compare fast sampling algorithms DDIM and CTD, along with our method RACTD on Maze2d environment. DDIM performs fast sampling based on a DDPM model, while CTD and RACTD (ours) distill an EDM teacher. The number of function evaluations (NFE) reflects the sampling speed of each algorithm. Results are reported as the mean and standard error over 100 random seeds.
Method NFE U-Maze Medium Large Average Score
Score Score Score
DDPM $100$ 126.3±$\pm 0.7$ $126.8$$\pm 3.0$ $144.8$$\pm 4.9$ $132.6$
DDIM 10 $121.2$ ±$\pm 1.1$ $126.2$$\pm 2.8$ $143.1$$\pm 4.9$ $130.2$
DDIM 1 $3.5$$\pm 4.7$ $-2.6$$\pm 12.5$ $-1.5$$\pm 0.7$ $-0.2$
EDM 80 $125.4$$\pm 0.6$ $120.1$$\pm 4.2$ 149.0 ±$\pm 0.5$ $131.5$
CTD 1 $123.4$$\pm 1.0$ $119.8$$\pm 4.1$ $127.1$$\pm 6.4$ $123.4$
RACTD (ours) 1 $125.7$$\pm 0.6$ 130.8 ±$\pm 1.8$ $143.8$$\pm 0.0$ 133.4
Appendix D The Trade-off between Mode Selection and Sample Diversity
In this section, we include a discussion about the trade-off between mode selection induced by our reward-aware training and the sample diversity of the student. Naturally, favoring selected modes can led to generation with limited sample diversity as summarized in Table 12. This trade-off between sample diversity and sample optimality observed in RACTD is similar to what has been seen in other generative domains (e.g., language model RLHF [Huang et al., 2024], classifier-guided diffusion, conditional image generation), where preference alignment also often reduces sample diversity. In our case, the reward model acts similarly to a classifier or an alignment reward model, guiding the model toward desirable behaviors and sacrificing some of the sample diversity by design.
Importantly, our decoupled framework allows the use of a single, unconditioned teacher with strong generalization capabilities across tasks. For multi-task or unseen-task settings, different reward models can be trained per task, and corresponding student models can be distilled from the same teacher using different reward models.
Table 12: A summarization of the trade off between sample diversity and model performance.
Sample diversity Performance Sample time
Reward agnostic diffusion High Low Slow
Reward aware diffusion Low High Slow
Reward agnostic consistency distillation High Low Fast
Reward aware consistency distillation Low High Fast
Appendix E Limitations and Future work
One limitation of our approach is the need to train three separate networks: the teacher, student, and reward model. Training the teacher can be time-consuming, as achieving strong performance often requires a higher number of denoising steps. Additionally, consistency trajectory distillation is prone to loss fluctuations, and incorporating a reward model into the distillation process may further amplify this instability. Future work will focus on developing a more stable and efficient training procedure, as well as exploring methods to integrate non-differentiable reward models into the framework.